### v1

@@
Abstract
@@
keywords:
Neural Networks, Activation function learning, Dynamic adjustment of structure
1 Introduction
In the field of artificial intelligence, scientists have been studying artificial neural networks almost as old as AI itself. Since the 1940s, the development of neural networks has had ups and downs. Since Warren McCulloch and Walter Pitts proposed the MP model, Frank Rosenblatt, Marvin Minsky Seymour Papert, Hinton, David Rumelhart, Sepp Hochreiter, Xavier Glorot, Yoshua Bengio, and a series of scientists in the process of the development of the neural network has played a very important role. All of them proposed or solved a series of important problems in the development of artificial neural networks. In today’s mainstream artificial neural network paradigm, once the structure is initialized, its activation function and network structure are fixed during the training process. This makes it difficult for the neural network to accurately adapt to the task. As a result, the network structure is either redundant or insufficient. Especially in the current context of various large models emerging endlessly and people blindly increasing the scale of neural networks in pursuit of more powerful performance. It is necessary to study a new neural network with dynamic learnable activation functions and adaptive adjustment of network structure. This is also a problem that future artificial intelligence research must face. Because human computing power is ultimately limited, the scale of the model cannot be expanded indefinitely.
Some traditional network structure optimization algorithms treat the network structure as the hyperparameters of neural networks and then use some optimization algorithms to optimize these hyperparameters. Such as evolutionary algorithms[1] and Bayesian optimization algorithms[2]. Recently, some automatic design methods of network structure have been proposed which are more flexible than the traditional hyperparameter optimization methods. Zoph and Le [3] define recurrent neural networks (RNNS) that generate a neural network architecture for a specific target problem, optimize the RNNS by using a strategy gradient approach, and find a good neural network architecture. The work of Real et al. 2017[4]; Suganuma, [5] uses evolutionary algorithms to optimize layer-to-layer neuronal connections to build high-performance neural network architectures. [6] consider a probability distribution that generates network structures, and optimizes the parameters of the distribution instead of directly optimizing the network structure. The Jagtap[7] paper presents a new type of neural network, Kronecker Neural Networks (Kronecker NNs), which forms a general framework for neural networks with adaptive activation functions. In addition, the paper proposes a Rowdy activation function, which aims to eliminate any saturated regions by injecting sinusoidal fluctuations, which include trainable parameters. Agostinelli[8] designed a new form of parametric piecewise linear activation functions that are learned independently for each neuron using gradient descent and can represent both convex and non-convex functions of the input.
We propose a new neural network called X-Net. Both the activation function and the network structure of X-Net are dynamically adaptive. While learning the type of activation function, the network structure can also be adjusted accurately at the neuron level. So that the neural network structure can automatically grow or shrink according to the demand to adapt to different difficult tasks. We design X-Net mainly from the following two aspects:
  • Increase the type of activation function to improve the unit representation ability of the network.
  • The network structure is dynamically adjusted in real-time, which makes the network structure match the task requirements accurately and reduces the occurrence of parameter redundancy and insufficiency.
First, we propose a new model that allows X-Net to theoretically use any differentiable function as the activation function, and the type of activation function can be selected in real-time according to the needs of the task under the guidance of the gradient during the network training. In particular, in the process of selecting the activation function, the precise adjustment of the network structure can be realized, so that the network structure and scale can be adapted to the current task as much as possible.
In summary, the X-Net model not only opens up a new research approach in the field of neural networks but also sets up a new technical standard for building truly adaptive intelligent systems. It is expected that X-Net can play its value in various fields in the future and promote the progress and development of mankind. In the end, we hope that our research will throw a brick to attract jade and stimulate enthusiasm for this new type of neural network.
Table 1: Comparison of extrapolation between X-Net, ANN, and SVR.
Benchmark X-Net ANN SVR Kronecker NN Shinichi[6] X-Net (No parameters)
In [-2,2] Not in [-2,2] In [-2,2] Not in [-2,2] In [-2,2] Not in [-2,2][-2,2] In[-2,2] Not in [-2,2] In [-2,2] Not in [-2,2] In [-2,2] Not in [-2,2]
Nguyen-1 $1.00_{\pm 0.0}$ $1.00_{\pm 0.00}$ $0.99_{\pm 0.08}$ $-29.34_{\pm 4.23}$ $0.99_{\pm 0.02}$ $-59.33_{\pm 1.21}$ $0.99_{\pm 0.05}$ $-32.45_{\pm 1.92}$ $0.99_{\pm 0.02}$ $-19.29_{\pm 0.84}$ $1.00_{\pm 0.00}$ $1.00_{\pm 0.00}$
Nguyen-2 $1.00_{\pm 0.00}$ $1.00_{\pm 0.00}$ $0.99_{\pm 0.04}$ $-111.57_{\pm 14.34}$ $0.99_{\pm 0.05}$ $-441.08_{\pm 3.53}$ $0.99_{\pm 0.03}$ $-167.28_{\pm 4.31}$ $0.99_{\pm 0.03}$ $-147.26_{\pm 2.01}$ $1.00_{\pm 0.00}$ $1.00_{\pm 0.00}$
Nguyen-3 $0.99_{\pm 0.06}$ $0.96_{\pm 0.04}$ $0.99_{\pm 0.05}$ $-553.86_{\pm 79.18}$ $0.99_{\pm 0.07}$ $-397_{\pm 04.22}$. $0.99_{\pm 0.02}$ $-504.81_{\pm 8.54}$ $0.99_{\pm 0.01}$ $-382.64_{\pm 6.30}$ $0.99_{\pm 0.02}$ $0.94_{\pm 0.03}$
Nguyen-4 $0.99_{\pm 0.09}$ $0.94_{\pm 0.06}$ $0.98_{\pm 0.10}$ $-2452.16_{\pm 282.23}$ $0.99_{\pm 0.08}$. $-5107_{\pm 32}$. $0.99_{\pm 0.03}$ $-3048.24_{\pm 202}$ $0.99_{\pm 0.02}$ $-3012.62_{\pm 26.44}$ $0.99_{\pm 0.03}$ $0.84_{\pm 0.09}$
Nguyen-5 $0.99_{\pm 0.04}$ $-0.12_{\pm 0.08}$ $0.99_{\pm 0.08}$ $-9.76_{\pm 0.82}$. $0.57_{\pm 0.07}$. $-19.81_{\pm 0.13}$ $0.99_{\pm 0.07}$ $-16.73_{\pm 0.93}$. $0.99_{\pm 0.04}$ $-14.28_{\pm 0.94}$ $0.86_{\pm 0.07}$ $0.10_{\pm 0.08}$
Nguyen-6 $1.00_{\pm 0.00}$ $1.00_{\pm 0.00}$ $0.99_{\pm 0.07}$ $-8.49_{\pm 1.30}$ $0.99_{\pm 0.05}$. $-2.34_{\pm 0.08}$. $0.99_{\pm 0.02}$ $-12.86_{\pm 0.79}$ $0.99_{\pm 0.04}$ $-13.72_{\pm 0.74}$ $1.00_{\pm 0.00}$ $1.00_{\pm 0.00}$
Nguyen-7 $0.99_{\pm 0.02}$ $0.94_{\pm 0.03}$ $0.99_{\pm 0.06}$ $0.62_{\pm 0.09}$. $0.99_{\pm 0.03}$ $-53.00_{\pm 0.46}$ $0.99_{\pm 0.03}$ $-0.29_{\pm 0.02}$ $0.99_{\pm 0.06}$ $0.12_{\pm 0.03}$ $0.99_{\pm 0.03}$ $0.62_{\pm 0.08}$
Nguyen-8 $1.00_{\pm 0.00}$ $1.00_{\pm 0.00}$ $0.99_{\pm 0.01}$ $0.71_{\pm 0.02}$. $0.96_{\pm 0.05}$. $-47.52_{\pm 0.20}$ $0.99_{\pm 0.03}$ $0.16_{\pm 0.01}$. $0.99_{\pm 0.03}$. $0.31_{\pm 0.01}$ $1.00_{\pm 0.00}$ $1.00_{\pm 0.00}$
Nguyen-9 $1.00_{\pm 0.00}$ $1.00_{\pm 0.00}$ $0.99_{\pm 0.05}$ $-3.31_{\pm 0.49}$. $0.98_{\pm 0.04}$. $-48.52_{\pm 0.19}$ $0.99_{\pm 0.06}$ $-10.17_{\pm 0.89}$. $0.99_{\pm 0.04}$ $-11.27_{\pm 1.20}$ $1.00_{\pm 0.00}$ $1.00_{\pm 0.00}$
Nguyen-10 $1.00_{\pm 0.00}$ $1.00_{\pm 0.00}$ $0.99_{\pm 0.09}$ $-15.56_{\pm 1.22}$ $0.98_{\pm 0.41}$ $-3.05_{\pm 0.41}$ $0.99_{\pm 0.05}$. $-28.75_{\pm 0.99}$ $0.99_{\pm 0.02}$ $-16.88_{\pm 0.96}$ $1.00_{\pm 0.00}$ $1.00_{\pm 0.00}$
Nguyen-11 $1.00_{\pm 0.00}$ $1.00_{\pm 0.00}$ $0.99_{\pm 0.02}$ $-124985_{\pm 4476.20}$ $0.99_{\pm 0.02}$ $-858911_{\pm 10006}$ $0.99_{\pm 0.1}$ $-204823_{\pm 3013}$ $0.99_{\pm 0.08}$ $-116424_{\pm 822}$ $1.00_{\pm 0.00}$ $1.00_{\pm 0.00}$
Nguyen-12 $0.99_{\pm 0.09}$ $-2.34_{\pm 0.10}$ $0.99_{\pm 0.20}$ $-599.34_{\pm 28.44}$ $0.99_{\pm 0.20}$ $-8466_{\pm 244.55}$ $0.99_{\pm 0.07}$ $-1027.13_{\pm 20.66}$ $0.99_{\pm 0.08}$ $-646.22_{\pm 12.64}$ $0.99_{\pm 0.09}$ $0.16_{\pm 0.04}$
Average 0.996 0.531 0.989 -10730.58 0.950 -72795 0.990 -17472 0.990 -10057 0.985 0.805
2 Results
@@
Evaluating X-Net. In order to comprehensively evaluate the performance of X-Net, we make a comprehensive evaluation of X-Net in terms of fitting ability and classification ability. In particular, we also test X-Net’s ability to scientific discovery on datasets from various disciplines such as environment and energy.
2.1 Fit ability evaluation
We compare X-Net with support vector regression(SVR), ANN, Kronecker NN, and Shinichi[6] four excellent regression algorithms.
  • Support Vector Regression (SVR). SVR is characterized by the use of the kernel, sparse solution, and VC to control the margin and the number of support vectors. One of the main advantages of SVR is that its computational complexity does not depend on the dimensions of the input space. In addition, it has excellent generalization ability and high prediction accuracy.
  • Artifical Neural Network (ANN). ANN is a kind of neural network which uses a backpropagation algorithm to update parameters. This network is one of the most widely used neural networks and is the foundation of deep learning. ANN has achieved excellent results in the field of regression.
  • Shinichi[6]. This method generates the probability distribution of the network structure and then optimizes the parameters of the distribution, rather than optimizing the network structure directly.
2.1.1 Analysis of fitting ability
@@
We tested the above four algorithms on the Nguyen data set. The dataset contains 12 curves, each of which is a mathematical formula. See the appendix for details. We used $R^{2}$ to test the fitting ability of our algorithm and SVR, ANN, Kronecker NN, and Shinichi[6]. The formula for $R^{2}$ is as follows(1).
@@
where $y_{i}$ is the true label value of the $i^{th}$sampling point, $\hat{y}_{i}$ is the value predicted by the model for the $i^{th}$ data, $\overline{y}$ is the mean of the true values $y$. The closer $R^{2}$ is to 1 the better the algorithm fits the target curve, and conversely, the farther $R^{2}$ is from 1, the worse the algorithm fits the target curve. The specific results are shown in Table 1[-2,2]. From the table, we can clearly find that our algorithm has better inherited the powerful nonlinear fitting ability of the neural network, and its fitting ability is not weaker than that of excellent regression algorithms such as SVR, ANN, Kronecker NN and Shinichi[6].
Table 2: A comparative analysis is presented, focusing on the accuracy and complexity of the final network structures across four algorithms in both regression and classification tasks.
REGRESSION
Benchmark X-Net ANN Kronecker NN Shinichi[6]
Acc Nod Para Acc Nod Para Acc Nod Para Acc Nod Para
Nguyen-1 1.0 5 22 0.99 14 78 0.99 12 61 0.99 12 52
Nguyen-2 1.0 9 38 0.99 18 118 0.99 16 97 0.99 15 79
Nguyen-3 1.0 14 58 0.99 18 118 0.99 14 78 0.99 11 47
Nguyen-4 0.99 20 82 0.98 28 253 0.99 24 193 0.99 18 115
Nguyen-5 0.99 5 16 0.99 26 222 0.99 28 253 0.99 30 289
Nguyen-6 0.99 5 18 0.99 8 33 0.99 8 33 0.99 7 27
Nguyen-7 1.0 5 16 0.99 12 61 0.99 10 46 0.99 8 37
Nguyen-8 0.99 1 4 0.99 16 97 0.99 16 97 0.99 12 61
Nguyen-9 0.99 3 10 0.99 40 481 0.99 28 253 0.99 21 139
Nguyen-10 0.99 4 14 0.99 130 4468 0.99 80 1761 0.99 56 919
Nguyen-11 0.99 1 6 0.99 16 97 0.99 20 141 0.99 12 65
Nguyen-12 1.0 9 40 0.99 16 97 0.99 12 61 0.99 14 73
Average 0.996 6.75 27 0.989 28.50 510.25 0.990 21.5 256.17 0.990 18 158.58
CLASSION
Iris 0.987 26 28 0.980 51 828 0.983 38 632 0.986 32 546
Mnist(6-dim) 0.894 55 166 0.886 201 11713 0.888 142 6735 0.889 128 6022
Mnist 0.995 316 4728 0.992 488 158260 0.992 463 117119 0.990 444 100626
(a) Fitting capability and extrapolation
(b) Performance of the Ada-$\alpha$ Activation Function
(c) Boston Housing Price Data Correlation Matrix.
(d) Predicting housing prices using the ’RM’
(e) Predicting housing prices using the ’LSTAT’
(f) Predicting housing prices using ’RM’ and ’LSTAT’

### v2

@@
A Novel Paradigm for Neural Computation: X-Net with Learnable Neurons and Adaptable Structure
Yanjie Li,13 Weijun Li,123∗ Lina Yu1, Min Wu1, Jingyi Liu1, Wenqiang Li13,
Linjun Sun1, Meilan Hao1, Shu Wei13, Yusong Deng13, LipingZhang1, XiaoliDong1,
HongQin1, XinNing1, YuguiZhang1, BaoliLu1, JianXu1, ShuangLi1
1AnnLab, Institute of Semiconductor, Chinese Academy of Sciences
HaiDian, Beijing, 100083, CN
2School of Integrated Circuits, University of Chinese Academy of Sciences
HuaiRou, Beijing, 101408, CN
3School of Electronic, Electrical and Communication Engineering
University of Chinese Academy of Sciences
HuaiRou, Beijing, 101408, CN
To whom correspondence should be addressed; E-mail: wjli@semi.ac.cn
@@ A Novel Paradigm for Neural Computation: X-Net with Learnable Neurons and Adaptable Structure
Introduction
Multi-Layer Perceptron (MLP) is the cornerstone of contemporary artificial intelligence. In the field of artificial intelligence, the study of MLP (?, ?, ?, ?, ?) is nearly as old as AI itself. However, there is an irreconcilable contradiction between the scale and performance of the MLP. To obtain better performance, it is often necessary to continuously expand the scale of the network in terms of depth and breadth. For example, the performance of GPT-1 to GPT-4 (?, ?, ?, ?, ?) is getting better and better, but the number of parameters rapidly increases from 1.3B to a trillion level, which brings the problem of energy consumption, calculation, storage, communication, and other costs (?). The high cost will hinder its application and promotion. So why do current neural networks tend to be so big? Its technical roots lie in two points:
1), The neuron activation function of classical neural networks is single and fixed, and its representation ability is relatively insufficient. Many neurons are often required to fit other types of nonlinear functions. Moreover, as the dimensionality of the problem increases, the number of neurons explodes exponentially.
2), The network structure as a hyperparameter needs to be predetermined and kept fixed during training. But in fact, it is difficult for us to rely on artificial experience to obtain an optimal network structure, and it is easy to have structural redundancy or insufficiency.
Figure 1: Figure illustrates the dynamic changes in the network structure of X-Net during the training process.
So, it is necessary to study a new generation of neural networks with dynamic learnable activation functions and adaptive adjustment of network structure. This is also a problem that future artificial intelligence research must face. Because human computing power is ultimately limited, the scale of the model cannot be expanded indefinitely.
In order to overcome the above problems, we try to explore a new generation of neural networks. We propose a new neural network called X-Net.
Table 1: The characteristics of X-Net and MLPs.
Neuron Type Neuron Learnability Weight Learnability Structure
MLPs Single × Pre-defined
X-Net Multiple Self-learnable
Table1 shows the characteristics of X-Net and MLP. The X-Net theoretically can use any differentiable function as the activation function, The experimental results show that the representation ability is greatly improved compared with MLP0.1. The type of activation function can be learned in real-time according to the needs of the task under the guidance of the gradient during the network training. In particular, The network structure is dynamically adjusted in real-time, which makes the network structure match the task requirements accurately and reduces the occurrence of redundancy and insufficiency. Fig. 1 shows a schematic diagram of the training process of X-Net.
In addition, we found that when the activation function uses some mathematical notation with nice properties (e.g., sin() has periodicity, etc.), X-Net can find some interpretable mathematical formulas in some cases. This will greatly improve the problem of uninterpretable results of traditional neural networks. It also shows that X-Net has great potential in facilitating scientific discovery. We believe this will be extremely attractive to researchers from other disciplines (biology, physics, materials, etc.) and we think it will drive the field of AI for Science to flourish even more.
In summary, the X-Net model not only opens up a new research direction in the field of neural networks but also sets up a new technical standard for building truly adaptive intelligent systems. In addition, X-Net has good universality and can empower the development of various disciplines. It has a very broad research prospect and research space. Finally, we hope that our study will provoke a reflection on the inadequacy of current neural network architectures, as well as generate enthusiasm for the exploration of next-generation neural networks.
Results
@@
In order to comprehensively evaluate the performance of X-Net, we make a comprehensive evaluation of X-Net on regression and classification tasks. In particular, we also test X-Net’s ability to make scientific discoveries on datasets from various disciplines such as environment and energy.
0.1 Representation capability with learnable activation functions
In order to compare the representation capabilities of X-Net and MLPs, we designed a unique experiment: we initialized the X-Net and MLP with three layers of neurons with 4, 2, and 1 neurons per layer. The difference is that the different layers in X-Net are sparsely connected in the form of a binary tree, whereas the MLP is fully connected. Each neuron in X-Net has a different activation function, whereas neurons in MLP all have a single ReLU activation function. We optimize X-Net and MLP using the LBFGS. The results in Table 2 show that X-Net exhibits a better fitting ability on the Nguyen dataset compared to MLPs, despite being sparsely connected. This result demonstrates that the X-Net can achieve superior nonlinear representation ability than the fully connected MLPs due to the diversity of activation functions. This also proves that the hypothesis of improving the representation ability of neural networks by increasing the type of activation function is feasible.
Table 2: The representation ability of using multiple activation function types versus a single activation function type for the same number of neurons.
Data Benchmark Activation Type
Multiple Single
Nguyen-1 $x_{1}^{3}+x_{1}^{2}+x_{1}$ 0.9999 0.5924
Nguyen-2 $x_{1}^{4}+x_{1}^{3}+x_{1}^{2}+x_{1}$ 0.9999 0.3922
Nguyen-3 $x_{1}^{5}+x_{1}^{4}+x_{1}^{3}+x_{1}^{2}+x_{1}$ 0.9999 0.7645
Nguyen-4 $x_{1}^{6}+x_{1}^{5}+x_{1}^{4}+x_{1}^{3}+x_{1}^{2}+x_{1}$ 0.9995 0.8023
Nguyen-5 $\sin(x_{1}^{2})\cos(x)-1$ 0.9956 0.2217
Nguyen-6 $\sin(x_{1})+\sin(x_{1}+x_{1}^{2})$ 0.9995 0.4337
Nguyen-7 $\log(x_{1}+1)+\log(x_{1}^{2}+1)$ 0.9999 0.6902
Nguyen-8 $\sqrt{x}$ 0.9999 0.6756
Nguyen-9 $\sin(x)+\sin(x_{2}^{2})$ 0.9940 0.7726
Nguyen-10 $2\sin(x)\cos(x_{2})$ 0.9859 0.8126
Nguyen-11 $x_{1}^{x_{2}}$ 0.9879 0.6673
Nguyen-12 $x_{1}^{4}-x_{1}^{3}+\frac{1}{2}x_{2}^{2}-x_{2}$ 0.9824 0.8615
Average 0.9954 0.6572
0.2 Complexity compared with MLPs
Artificial neural networks have a fixed network structure and a number of nodes, which can easily lead to redundancy in parameters and nodes, greatly slowing down the convergence efficiency of the neural network. In contrast, X-Net is more flexible, with a dynamically changing network structure and the number of nodes that can be adaptively adjusted based on the complexity of the problem. This greatly alleviates the problem of redundancy in nodes and parameters in ordinary neural networks. To test the complexity of the model obtained by four algorithms X-Net, and MLP on the same task, we stop the training when $R^{2}=0.99$, and count the number of nodes and parameters used by the two networks at this time. The specific statistical results are shown in Table 3. From Table 3(REGRESSION), it is evident that in fitting the Nguyen data set, the MLP necessitates approximately fourfold the node count in comparison to X-Net and demands almost a twentyfold increase in the number of parameters for optimization. The results in Table 3(CLASSION) show that X-Net requires an average number of nodes comparable to that of MLP for the classification task, but the number of weight parameters is only about 1.4% of that of MLP, and under these conditions, X-Net exhibits performance comparable to that of MLP.
0.3 Performance on regression tasks
0.3.1 Fitting accuracy
Table 3: Comparison of various indicators between X-Net and MLP on regression and Classion tasks
REGRESSION
Data Benchmark X-Net MLP
Fitting($R^{2}$) Nodes Parameters Fitting($R^{2}$) Nodes Parameters
Nguyen-1 $x_{1}^{3}+x_{1}^{2}+x_{1}$ $1.0000_{\pm 0.00}$ 5 12 $0.9999_{\pm 0.08}$ 14 78
Nguyen-2 $x_{1}^{4}+x_{1}^{3}+x_{1}^{2}+x_{1}$ $1.0000_{\pm 0.00}$ 9 38 $0.9999_{\pm 0.04}$ 18 118
Nguyen-3 $x_{1}^{5}+x_{1}^{4}+x_{1}^{3}+x_{1}^{2}+x_{1}$ $0.9999_{\pm 0.06}$ 14 58 $0.9999_{\pm 0.05}$ 18 118
Nguyen-4 $x_{1}^{6}+x_{1}^{5}+x_{1}^{4}+x_{1}^{3}+x_{1}^{2}+x_{1}$ $0.9999_{\pm 0.09}$ 20 82 $0.9998_{\pm 0.10}$ 28 253
Nguyen-5 $\sin(x_{1}^{2})\cos(x)-1$ $0.9998_{\pm 0.04}$ 5 16 $0.9984_{\pm 0.08}$ 26 222
Nguyen-6 $\sin(x_{1})+\sin(x_{1}+x_{1}^{2})$ $1.0000_{\pm 0.00}$ 6 18 $0.9996_{\pm 0.07}$ 8 333
Nguyen-7 $\log(x_{1}+1)+\log(x_{1}^{2}+1)$ $0.9996_{\pm 0.02}$ 5 16 $0.9998_{\pm 0.06}$ 12 61
Nguyen-8 $\sqrt{x}$ $1.0000_{\pm 0.00}$ 1 4 $0.9999_{\pm 0.01}$ 16 97
Nguyen-9 $\sin(x)+\sin(x_{2}^{2})$ $1.0000_{\pm 0.00}$ 4 14 $0.9984_{\pm 0.05}$ 40 481
Nguyen-10 $2\sin(x)\cos(x_{2})$ $1.0000_{\pm 0.00}$ 7 24 $0.9994_{\pm 0.09}$ 130 4468
Nguyen-11 $x_{1}^{x_{2}}$ $1.0000_{\pm 0.00}$ 3 10 $0.9999_{\pm 0.02}$ 16 97
Nguyen-12 $x_{1}^{4}-x_{1}^{3}+\frac{1}{2}x_{2}^{2}-x_{2}$ $0.9996_{\pm 0.09}$ 9 40 $0.9988_{\pm 0.20}$ 16 97
Average 0.9999 7.33 28.50 0.9995 28.50 510.25
CLASSION
Data Benchmark X-Net MLP
Accuracy Nodes Parameters Accuracy Nodes Parameters
Dataset-1 Iris 98.7% 28 112 99.0% 67 1315
Dataset-1 Mnist(6-dim) 89.4% 65 244 88.6% 298 23342
Dataset-1 Mnist 99.5% 816 3084 99.2% 788 267612
Dataset-1 Fashion-MNIST(6-dim) 76.2% 122 486 75.1% 398 31143
Dataset-1 Fashion-MNIST 94.1% 1066 3884 94.4% 1244 544129
Dataset-1 CIFAR-10(6-dim) 26.4% 206 764 24.6% 414 35292
Dataset-1 CIFAR-10 46.8% 2733 10072 48.4% 2164 876503
Average 75.87% 719.43 2663.71 75.61% 767.57 234190.86
@@ 0.3.1 Fitting accuracy
We tested the above four algorithms on the Nguyen data set. The dataset contains 12 curves, each of which is a mathematical formula. See the appendix for details. We used $R^{2}$ to test the fitting ability of our algorithm and MLP. The formula for $R^{2}$ is as follows(1).
@@
where $y_{i}$ is the true label value of the $i^{th}$sampling point, $\hat{y}_{i}$ is the value predicted by the model for the $i^{th}$ data, $\overline{y}$ is the mean of the true values $y$. The closer $R^{2}$ is to 1 the better the algorithm fits the target curve, and conversely, the farther $R^{2}$ is from 1, the worse the algorithm fits the target curve. The specific results are shown in Table 3 (REGRESSION). From the table, we can clearly find that our algorithm has better inherited the powerful nonlinear fitting ability of the neural network, and its fitting ability is not weaker than that of excellent regression algorithms such as MLP.
0.4 Performance on classification tasks
The classification task stands as a quintessential sub-domain within the realm of machine learning. Consequently, gauging the classification performance of algorithms becomes imperative. To validate the classification efficacy of X-Net, we employed the Iris dataset, MNIST, Fishion-MNIST, and CIFAR-10 dataset as our experimental datasets. We juxtaposed the performance of X-Net with the MLP. On the MNIST, Fishion-MNIST, and CIFAR-10 datasets, we conducted two distinct experiments: initially, we employed PCA to dimensionally reduce the data, transmuting the images from a p dimension(Number of pixels) down to 6 dimensions, and subsequently classifying using the reduced data. The secondary experiment involved direct classification without any dimension reduction.
Table 3 (CLAASSION) meticulously delineates the performance comparison between X-Net and MLP on the Iris, MNIST, Fashion-MNIST, and CIFAR-10 datasets. The outcomes suggest that X-Net marginally outperforms the conventional MLP across all seven classification tasks. Regarding network structural complexity, X-Net surpasses its counterpart. Taking the Iris dataset as a case in point, the neuron count of X-Net is merely half that of MLP, with its parameter magnitude being a mere tenth of that of MLP. On the MNIST dataset, In the experimentation employing PCA, the node count of X-Net is but a quarter of MLP, and its parameter magnitude stands at a mere $1.04\%$ (1/96) of MLP. In experiments without dimension reduction, the node count of X-Net amounts to two-thirds that of MLP, while the parameter count is a mere $1.1\%$ (1/90) of MLP. The same trend is also shown in Fishion-MNIST and CIFAR-10.
All in all, X-Net achieves comparable accuracy to MLP on classification tasks; however, its network structure complexity is much lower than MLP.
0.5 Used in scientific discovery
0.5.1 Modeling in space science (Airfoil-self-noise) (?)
(a) Results of fitting the Airfoil-self-noise data (Eq.5)
(b) Figure a after sorting by $y_{pre}$
(c) Correlation coefficient matrix for Airfoil-self-noise
(d) Results of fitting the Airfoil-self-noise data(Eq.6)
(e) Figure d after sorting by $y_{pre}$
(f) Global temperature change prediction outcomes